a rectangle has a lenght that is 3 yard greater than the width.
the area is 180 find lengs and width?

Answer by stanbon(75887)

a rectangle has a length that is 3 yard greater than the width.
the area is 180 find length and width?
------------
Let width be "x" yrds.
Then length is "x+3" yrds.
--------------------
EQUATION:
x(x+3) = 180
x^2+3x-180 = 0
(x-12)(a+15) = 0
Positive solution:
x=12 yrds (width)
x+3 = 15 yrds (length)
=======================
Cheers,
Stan H.